OpenStudy (anonymous):

WILL FANN & MEDALL In a word game you choose a tile from a bag, replace it, and then choose another. If there are 14 vowels and 22 constants, what is the probability you will choose a constant and then a vowel? a. 77/324 b. 11/45 c. 45/11 d. 2/9

OpenStudy (tkhunny):

What is p(consonant on the 1st draw)?

OpenStudy (sparrow2):

a

OpenStudy (sparrow2):

22/36(constants) x 14/36(vowel) = 11/18 x 7/18=77/324

OpenStudy (anonymous):

wow!! thxx :)

OpenStudy (tkhunny):

I'm going to guess that should be "consonants", not "constants".

OpenStudy (anonymous):

nope... its "constants"

OpenStudy (tkhunny):

Weird. If only that meant anything. In any case, there are 36 total tiles (14+22). p(grabbing a consonant) is 22/36 Since you put it back, the next draw in totally independent. There are still 36 tiles but only 14 that will lead to success as it is defined.

OpenStudy (anonymous):

confused .___. so is a right or wrong?

OpenStudy (tkhunny):

You have enough information. What is YOUR solution?

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

is it right........

OpenStudy (tkhunny):

That's what sparrow said. What is YOUR solution. There are only four possible outcomes. Consonant, Consonant Consonant, Vowel Vowel, Consonant Vowel, Vowel What are the probabilities of each? (22/36), (22/36) (22/36), (14/36) (14/36), (22/36) (14/36), (14/36) Which is it asking?

OpenStudy (anonymous):

its asking for what is the probability you will choose a constant then a vowel..

OpenStudy (tkhunny):

Perfect. Which of the four I laid out is that?

OpenStudy (anonymous):

(22/36), (14/36)

OpenStudy (anonymous):

right?

OpenStudy (tkhunny):

Very good. Note: Part of the study of mathematics is learning how to think. Learning how to put together the story. Learning how to piece together the puzzle. The correct answer is good, but that is just not the reason why you are in a math class - why your school makes you take a class like this. Figure out the rest. I have little doubt that you can do it. You just have to decide that it may require a kind of thinking you are not used to.

OpenStudy (anonymous):

okay..thxx.. after i get the answer can i ask u if its right?

OpenStudy (tkhunny):

My point is - YOU need to know it is right. You need to tell me! :-). You have it right, already. The rest is arithmetic.

OpenStudy (anonymous):

oh okay.... thxx!!!! :)